If f(x)=4x-3 find the following:
f(-1)
McGraw Hill 6th edtion, chapter 7.5 exercise #30
Thanks for your time!!!!

Answer by stanbon(48558)

If f(x)=4x-3 find the following:
f(-1)=4(-1)-3=-4-3=-7
Cheers,
Stan H.